# Polynomial-Time Preparation of Low-Temperature Gibbs States for 2D Toric Code

### v1

@@ 4.2.1 Analysis of the quasi-1D structure
$\mathrm{Gap}\left(-\mathcal{L}_{\rm local}\middle|_{\mathcal{B}_{{\bf 1},{\bf 1}}}\right)=\max\left\{e^{-\mathcal{O}(\beta)},\Omega(N^{-3})\right\}\,.$
@@
$\mathrm{Gap}\left(-\mathcal{L}_{1}\middle|_{\mathcal{A}^{\rm full}_{\rm m}}\right)=\max\left\{e^{-\mathcal{O}(\beta)},\Omega(N^{-3})\right\}\,.$ (4.33)
@@
$\mathrm{Gap}\left(-\mathcal{L}_{2}\middle|_{\mathcal{A}^{\rm full}_{\rm e}}\right)=\max\left\{e^{-\mathcal{O}(\beta)},\Omega(N^{-2.5})\right\}\,.$
@@
We next decompose the space $\mathcal{A}^{\rm full}_{\rm m}$ such that $\mathcal{L}_{1}$ presents a block diagonal form. Note that $\mathcal{A}^{\rm full}_{\rm m}$ is spanned by |$\ket{m^{\prime}}\bra{m}$ | with |$\ket{m},\ket{m^{\prime}}\in\mathcal{H}^{\rm m}_{\rm b}$ having an even number of -- signs. For any $\Lambda\subset\{1,2,\ldots,L^{2}\}$ with even $L^{2}-|\Lambda|$, we introduce the subspace $\mathcal{A}^{\rm full}_{\rm m}(\Lambda)$ of $\mathcal{A}^{\rm full}_{\rm m}$ spanned by |$\ket{m^{\prime}}\bra{m}$ |, where $m=m^{\prime}$ on $\Lambda$ and $m=-m^{\prime}$ on the the complement $\Lambda^{c}:=\{1,2,\ldots,L^{2}\}\backslash\Lambda$. It is straightforward to check that the subspaces $\mathcal{A}^{\rm full}_{\rm m}(\Lambda)$ are orthogonal with respect to both the GNS and HS inner products, and thereby
@@
  • If $\Gamma_{\rm ab}=\{1,2,\ldots,L^{2}\}$ (i.e., $\Lambda=[L^{2}]$), [2, Proposition 2] has given a spectral gap lower bound that exponentially decays in $\beta$. However, this lower bound is far from sharp at low temperature. Next, we shall derive a lower bound of gap polynomially decaying in $N$ but independent of $\beta$, which is summarized in the following lemma.
@@
    $\mathrm{Gap}\left(-\mathcal{L}_{1}\middle|_{\mathcal{A}^{\rm full}_{\rm m}\left([L^{2}]\right)}\right)=\Omega(N^{-3})\,.$
Once Lemma 13 is proved, we can combine it with Lemma 12, as well as (4.36) and (4.38), to obtain the desired (4.33):
$\mathrm{Gap}\left(-\mathcal{L}_{1}|_{\mathcal{A}_{\rm m}^{\rm full}}\right)\geq\min\left\{\frac{1}{2},\mathrm{Gap}\left(-\mathcal{L}_{1}\middle|_{\mathcal{A}^{\rm full}_{\rm m}\left([L^{2}]\right)}\right)\right\}=\Omega(N^{-3})\,.$
We next prove Lemma 13 to conclude the proof of Lemma 10. From (4.52), the matrix representation of the master Hamiltonian of $-\mathcal{L}_{1}$ can be written as:
$K_{L^{2}}^{\beta}=K_{L^{2}}+(K_{L^{2}}^{\beta}-K_{L^{2}})\,,$
with
and $K_{L^{2}}^{\beta}-K_{L^{2}}$ is given by
$\left\lVert K_{L^{2}}^{\beta}-K_{L^{2}}\right\rVert=\Theta(L^{2}e^{-2\beta})=\mathcal{O}(L^{-8})\,,\quad\text{when $\beta\geq 5\ln L$\,.}$
Therefore, to obtain the desired gap estimate (A.9), we only need to consider the gap at zero temperature and prove $\mathrm{Gap}(K_{L^{2}})\geq\Omega(L^{-6})$.
For this, we consider the following configuration space
[…]
$K_{L^{2}}:\mathcal{S}_{L^{2},k}\to\mathcal{S}_{L^{2},k}\,,$ (4.42)
$K_{L^{2}}=K_{L^{2}-1}\otimes I_{1}+I_{L^{2}-2}\otimes(k_{T}+k_{D})\,.$
In addition, there holds
$a^{*}K_{L^{2}}a$ $\geq a^{*}(K_{L^{2}-1}\otimes I_{1})a$ (4.43)
$=(a_{+})^{*}K_{L^{2}-1}a_{+}+(a_{-})^{*}K_{L^{2}-1}a_{-}$
$\geq\min\left\{\lambda_{L^{2}-1,k},\lambda_{L^{2}-1,k-1}\right\}\,,$
$K_{L^{2}}|_{\mathcal{A}_{L^{2},2}}=(K_{T})|_{\mathcal{A}_{L^{2},2}}+(K_{D})|_{\mathcal{A}_{L^{2},2}}\,,$
and $K_{T}(\varphi)=0$ for $\varphi\in\mathcal{A}_{L^{2},2}$ if and only if $I_{i}\otimes k_{T}\otimes I_{L^{2}-2-i}(\varphi)=0$ for all $i$, which implies
$\mathrm{Ker}\left(K_{T}|_{\mathcal{A}_{L^{2},2}}\right)=\mathrm{Span}\bigg{\{}\sum_{a\in\mathcal{A}_{L^{2},2}}a\bigg{\}}\,.$
Then, it follows that
$\bigg{(}\sqrt{\frac{2}{L^{2}(L^{2}-1)}}\sum_{a\in\mathcal{A}_{L^{2},2}}a\bigg{)}^{*}K_{D}\bigg{(}\sqrt{\frac{2}{L^{2}(L^{2}-1)}}\sum_{a\in\mathcal{A}_{L^{2},2}}a\bigg{)}\\ =\frac{4}{L^{2}(L^{2}-1)}\bigg{(}\sum_{a\in\mathcal{A}^{--}_{L^{2},2}}a\bigg{)}^{*}\bigg{(}\sum_{a\in\mathcal{A}^{--}_{L^{2},2}}a\bigg{)}=\frac{4}{L^{2}}\,,$
where $\mathcal{A}_{{L^{2}},2}^{--}:=\left\{a\in\left\{+,-\right\}^{\otimes{L^{2}}}\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ \text{$a$ has $2$ consecutive $``-"$ signs}\right\}$. By Lemma 3 (item 4) and $\|K_{D}|_{\mathcal{A}_{L^{2},2}}\|=\mathcal{O}(1)$, we have
$K_{T}|_{\mathcal{A}_{L^{2},2}}+K_{D}|_{\mathcal{A}_{L^{2},2}}\succeq\frac{(4/L^{2})\mathrm{Gap}\left(K_{T}|_{\mathcal{A}_{L^{2},2}}\right)}{\mathrm{Gap}\left(K_{T}|_{\mathcal{A}_{L^{2},2}}\right)+\mathcal{O}(1)}\,.$ (4.44)
We next estimate the spectral gap of $(K_{T})|_{\mathcal{A}_{L^{2},2}}$. We note from (4.39)-(4.40) that $K_{T}$ can be identified as a ferromagnetic Heisenberg-$1/2$ model on $L^{2}$ particles:
$K_{T}=-\frac{1}{2}\sum_{j=1}^{L^{2}-1}\left(\sigma_{j}^{x}\sigma_{j+1}^{x}+\sigma_{j}^{y}\sigma_{j+1}^{y}+\sigma_{j}^{z}\sigma_{j+1}^{z}-I\right)\,.$
From [35, Proposition 2] for the Heisenberg model, we have
$\mathrm{Gap}\left((K_{T})|_{\mathcal{A}_{L^{2},2}}\right)\geq\mathrm{Gap}\left(K_{T}\right)=\Omega(L^{-4})\,.$
Plugging this back into (4.44), we have
$\lambda_{L^{2},2}=\Omega\left(L^{-6}\right)\,.$ (4.45)
Using Eq. 4.43, it follows that $\min_{3\leq k\leq L^{2}}\lambda_{L^{2},k}=\Omega\left(L^{-6}\right)$ and $\mathrm{Gap}\left(-\mathcal{L}_{1}|_{\mathcal{A}^{\rm full}_{\rm m}([L^{2}])}\right)=\Omega\left(L^{-6}\right)$. To summarize, there holds
$\mathrm{Gap}\left(-\mathcal{L}_{1}\right)\geq\min\left\{\frac{1}{2},\mathrm{Gap}\left(-\mathcal{L}_{1}|_{\mathcal{A}^{\rm full}_{\rm m}([L^{2}])}\right)\right\}=\Omega\left(L^{-6}\right)=\Omega\left(N^{-3}\right)\,,$ (4.46)
for $\beta\geq 5\ln L$. This concludes the proof of Lemma 13, and consequently Lemma 10. ∎
@@
$\mathrm{Gap}\left(-\mathcal{L}_{2}\middle|_{\mathcal{A}^{\rm full}_{\rm m}\left([L^{2}]\right)}\right)=\Omega(N^{-2.5})\,.$
$\left\lVert K_{L^{2}}^{\beta}-K_{L^{2}}\right\rVert=\Theta(L^{2}e^{-2\beta})=\mathcal{O}(L^{-8})\,,\quad\text{for $\beta\geq 5\ln L$}\,,$
with
Then, $K_{L^{2}}$ is block diagonal for (4.47) and
We first consider the subspace $\mathcal{S}_{{L^{2}},2}$, whose basis vectors contain only two ``"``"``-"` ` - " signs. Our strategy for lower bounding $\lambda_{\min}\left(K_{L^{2}}|_{\mathcal{S}_{{L^{2}},2}}\right)$ is to reduce this problem to a straight line case as in Fig. 4. For this, we introduce a set of lines that covering all vertices in the comb. More specifically, we may regard the comb as a connected graph (more precisely, a tree) with $L^{2}$ vertices of degree at most 3, each corresponding to a star $\textbf{X}_{s}$ that interacts the comb. Let ${\bf D}_{\rm comb}$ be the set of degree-one vertices (i.e., end-points) in the comb, and $l_{u,v}$ be the shortest path between the vertices $u$ and $v$ in the comb. Define
For an arbitrary unit vector $\alpha\in\mathcal{S}_{{L^{2}},2}$:
$\alpha=\sum_{a\in\mathcal{A}_{L^{2},2}}p_{a}\ket{a}\,,$
by 16, there exists a path $\tilde{l}$ in ${\bf P_{\rm comb}}$ such that for
it holds that
$\|\alpha_{\tilde{l}}\|_{2}^{2}=\Omega(1/L^{2})\,.$ (4.49)
$(K_{L^{2}})_{\tilde{l}}:=\sum_{e=(u,v)\in\tilde{l}}(k_{T}+k_{D})_{u,v}\,,$
$=$ $\leavevmode\nobreak\ \Omega(L^{-2})\cdot\Omega(L^{-3})=\Omega(L^{-5})\,,$
where the first step follows from $K_{L^{2}}-(K_{L^{2}})_{\tilde{l}}$ is positive semi-definite, the second step follows from $\langle\alpha-\alpha_{\tilde{l}},(K_{L^{2}})_{\tilde{l}}\alpha_{\tilde{l}}\rangle=0$, the third step follows from $(K_{L^{2}})_{\tilde{l}}$ is equivalent to a 1D chain of length $|\tilde{l}|\leq\ell_{\rm comb}=O(L)$ and $a_{\tilde{l}}\in\mathcal{S}_{|\tilde{l}|,2}$, and the last step follows from (4.49) and (4.45).
Thus, we conclude
$K_{L^{2}}|_{\mathcal{S}_{{L^{2}},2}}=\Omega(L^{-5})=\Omega(N^{-2.5})\,.$
$\alpha=\sum_{a\in\mathcal{A}_{L^{2},k}}p_{a}\ket{a}\,,$
by 16, there exists a path $\tilde{l}$ in ${\bf P_{\rm comb}}$ such that for
we have
$\sum_{q=2}^{k}\|\alpha_{\tilde{l},q}\|_{2}^{2}=\Omega(1/L^{2})\,.$ (4.51)
$=$ $\leavevmode\nobreak\ \Omega(L^{-2})\cdot\Omega(L^{-3})=\Omega(L^{-5})\,,$
where the first step follows from $K_{L^{2}}-(K_{L^{2}})_{\tilde{l}}$ is positive semi-definite, the second step follows from $\alpha_{\tilde{l},q_{1}}^{*}(K_{L^{2}})_{\tilde{l}}\alpha_{\tilde{l},q_{2}}=0$ for $q_{1}\neq q_{2}$, the third step follows from $(K_{L^{2}})_{\tilde{l}}$ is equivalent to a 1D chain of length $|\tilde{l}|\leq\ell_{\rm comb}=O(L)$, and $\alpha_{\tilde{l},q}$ is in the subspace $\mathcal{S}_{|\tilde{l}|,q}$, the fourth step follows from the recursion relation (4.43) that $\lambda_{\ell_{\rm comb},q}\geq\lambda_{\ell_{\rm comb},2}$, an the last step follow from (4.45) and (4.51). Thus, we have
$K_{L^{2}}|_{\mathcal{S}_{{L^{2}},k}}=\Omega(L^{-5})=\Omega(N^{-2.5})\,.$
Combining them all together, we now can conclude
$\mathrm{Gap}\left(-\mathcal{L}_{2}|_{\mathcal{A}_{\rm e}^{\rm full}([L^{2}])}\right)\geq\min_{2\leq k\leq L^{2}}\lambda_{\min}\left(K_{L^{2}}|_{\mathcal{S}_{{L^{2}},k}}\right)-\mathcal{O}(N^{-4})=\Omega(N^{-2.5})\,.$
This concludes the proof of Lemma 15, and thus the proof of Lemma 11. ∎
Fact 16.
$\alpha=\sum_{a\in\mathcal{A}_{L^{2},k}}p_{a}\ket{a}\,,$
$\sum_{\begin{subarray}{c}a\in\mathcal{A}_{L^{2},k}:\\ \#\{s\in\tilde{l}:a_{s}=``-"\}>1\end{subarray}}|p_{a}|^{2}=\Omega(1/L)\,.$
Proof.
$=$ $\leavevmode\nobreak\ 1\,,$
$\sum_{\begin{subarray}{c}a\in\mathcal{A}_{L^{2},k}:\\ \#\{s\in l:a_{s}=``-"\}>1\end{subarray}}|p_{a}|^{2}\geq 1/L\,,$
which proves the proposition. ∎
@@
Figure 6: Examples of lines on the graph.
@@ 4.3 Proof of Lemmas
The decomposition (4.11) is straightforward by the construction. Here we only prove the fact that the algebra $\mathcal{A}^{\rm full}_{\rm m}=\mathcal{B}(\mathcal{H}_{\rm b}^{\rm m})$ can be generated by $\{\mathbf{Z}_{p}\}_{p}\cup\{\sigma^{x}_{j}\}_{j\in\text{\rm snake}}$. The claim for $\mathcal{A}^{\rm full}_{\rm e}$ can be similarly proved. Indeed, each basis (admissible bond) |$\ket{m}$ in $\mathcal{H}_{\rm b}^{\rm m}$ can be identified as a Pauli string $\sigma^{x}_{j_{1}}\cdots\sigma^{x}_{j_{k}}$ with $j_{i}$ on the snake. Each operator on $\mathcal{H}_{\rm b}^{\rm m}$ can be written as the linear combination of |$\ket{m^{\prime}}\bra{m}$ |, where |$\ket{m}$ and |$\ket{m^{\prime}}$ are admissible bonds. Meanwhile, each |$\ket{m^{\prime}}\bra{m}$ | with |$\ket{m^{\prime}}\neq\ket{m}$ is a composition of flips of neighboring states. Thus, it suffices to consider the case |$\ket{m^{\prime}}\bra{m}$ | that maps the state |$\ket{m}=\sigma^{x}_{j_{1}}\cdots\sigma^{x}_{j_{k}}\ket{+1^{L^{2}}}$ [^3] More rigorously, |$\ket{m}=\sigma^{x}_{j_{1}}\cdots\sigma^{x}_{j_{k}}\ket{+1^{L^{2}}}$ is defined as the bond configuration for $\{\textbf{Z}_{p}\}_{p}$ associated with the spin configuration $\sigma^{x}_{j_{1}}\cdots\sigma^{x}_{j_{k}}\ket{\psi_{o}}$, where |$\ket{\psi_{o}}$ is a ground state (4.5). to a neighboring one |$\ket{m^{\prime}}=\sigma^{x}_{j_{2}}\cdots\sigma^{x}_{j_{k}}\ket{+1^{L^{2}}}$, which can be easily constructed using $\{\textbf{Z}_{p}\}_{p}$ and $\{\sigma^{x}_{j}\}_{j\in{\rm snake}}$. Here we give the representation for |$\ket{m^{\prime}}\bra{m}$ |:
@@
The case of |$\ket{m}\bra{m}$ | can be similarly done. ∎

### v2

@@ 4.2.1 Analysis of the quasi-1D structure
$\mathrm{Gap}\left(-\mathcal{L}_{\rm local}\middle|_{\mathcal{B}_{{\bf 1},{\bf 1}}}\right)=\max\left\{e^{-\mathcal{O}(\beta)},\Omega(N^{-2})\right\}\,.$
@@
$\mathrm{Gap}\left(-\mathcal{L}_{1}\middle|_{\mathcal{A}^{\rm full}_{\rm m}}\right)=\max\left\{e^{-\mathcal{O}(\beta)},\Omega(N^{-2})\right\}\,.$ (4.33)
@@
$\mathrm{Gap}\left(-\mathcal{L}_{2}\middle|_{\mathcal{A}^{\rm full}_{\rm e}}\right)=\max\left\{e^{-\mathcal{O}(\beta)},\Omega(N^{-2})\right\}\,.$
@@
We next decompose the space $\mathcal{A}^{\rm full}_{\rm m}$ such that $\mathcal{L}_{1}$ presents a block diagonal form. Note that $\mathcal{A}^{\rm full}_{\rm m}$ is spanned by |$\ket{m^{\prime}}\bra{m}$ | with |$\ket{m},\ket{m^{\prime}}\in\mathcal{H}^{\rm m}_{\rm b}$ having an even number of -- signs. For any $\Lambda\subset\{1,2,\ldots,L^{2}\}$ with even $L^{2}-|\Lambda|$, we introduce the subspace $\mathcal{A}^{\rm full}_{\rm m}(\Lambda)$ of $\mathcal{A}^{\rm full}_{\rm m}$ spanned by |$\ket{m^{\prime}}\bra{m}$ |, where $m=m^{\prime}$ on $\Lambda$ and $m=-m^{\prime}$ on the complement $\Lambda^{c}:=\{1,2,\ldots,L^{2}\}\backslash\Lambda$. It is straightforward to check that the subspaces $\mathcal{A}^{\rm full}_{\rm m}(\Lambda)$ are orthogonal with respect to both the GNS and HS inner products, and thereby
@@
  • If $\Gamma_{\rm ab}=\{1,2,\ldots,L^{2}\}$ (i.e., $\Lambda=[L^{2}]$), [2, Proposition 2] has given a spectral gap lower bound that exponentially decays in $\beta$. However, this lower bound is far from sharp at low temperatures. Next, we shall derive a lower bound of gap polynomially decaying in $N$ but independent of $\beta$, which is summarized in the following lemma.
@@
    $\mathrm{Gap}\left(-\mathcal{L}_{1}\middle|_{\mathcal{A}^{\rm full}_{\rm m}\left([L^{2}]\right)}\right)=\Omega(N^{-2})\,.$
The proof of Lemma 13 is postponed to Section 4.3 for readability, which relies on a novel estimate of the minimal eigenvalue of a perturbed graph Laplacian. Once Lemma 13 is proved, we can combine it with Lemma 12, as well as (4.36) and (4.38), to obtain the desired (4.33):
$\mathrm{Gap}\left(-\mathcal{L}_{1}|_{\mathcal{A}_{\rm m}^{\rm full}}\right)\geq\min\left\{\frac{1}{2},\mathrm{Gap}\left(-\mathcal{L}_{1}\middle|_{\mathcal{A}^{\rm full}_{\rm m}\left([L^{2}]\right)}\right)\right\}=\Omega(N^{-2})\,.$
@@
$\mathrm{Gap}\left(-\mathcal{L}_{2}\middle|_{\mathcal{A}^{\rm full}_{\rm m}\left([L^{2}]\right)}\right)=\Omega(N^{-2})\,.$
The proof of Lemma 15 is deferred to Section 4.3. ∎
@@
Figure 5: Examples of lines on the graph. Both lines are the shortest path between two degree-one vertices (i.e., end-points) on the comb-like graph. The detailed application of these lines can be found in the proof of Lemma 15.
@@ 4.3 Proof of Lemmas
The decomposition (4.11) is straightforward by the construction. Here we only prove the fact that the algebra $\mathcal{A}^{\rm full}_{\rm m}=\mathcal{B}(\mathcal{H}_{\rm b}^{\rm m})$ can be generated by $\{\mathbf{Z}_{p}\}_{p}\cup\{\sigma^{x}_{j}\}_{j\in\text{\rm snake}}$. The claim for $\mathcal{A}^{\rm full}_{\rm e}$ can be similarly proved. Indeed, each basis (admissible bond) |$\ket{m}$ in $\mathcal{H}_{\rm b}^{\rm m}$ can be identified as a Pauli string $\sigma^{x}_{j_{1}}\cdots\sigma^{x}_{j_{k}}$ with $j_{i}$ on the snake. There are different ways to index the plaquette: and spins: One example is to index $\{\textbf{Z}_{p}\}_{p}$ as $\{\textbf{Z}_{j}\}_{j=1}^{L^{2}}$ along the snake from left to right and up to down, and index the spins from $2$ to $L^{2}$ on the snake in the same way so that each spin $j$ corresponds a pair of plaquettes $j-1$ and $j$, see Fig. 3 for detail. Each operator on $\mathcal{H}_{\rm b}^{\rm m}$ can be written as a linear combination of |$\ket{m^{\prime}}\bra{m}$ |, where |$\ket{m}$ and |$\ket{m^{\prime}}$ are admissible bonds. Meanwhile, each |$\ket{m^{\prime}}\bra{m}$ | with |$\ket{m^{\prime}}\neq\ket{m}$ is a composition of flips of neighboring states. Thus, it suffices to consider the case |$\ket{m^{\prime}}\bra{m}$ | that maps the state |$\ket{m}=\sigma^{x}_{j_{1}}\cdots\sigma^{x}_{j_{k}}\ket{+1^{L^{2}}}$ to a neighboring one |$\ket{m^{\prime}}=\sigma^{x}_{j_{2}}\cdots\sigma^{x}_{j_{k}}\ket{+1^{L^{2}}}$, which can be constructed using $\{\textbf{Z}_{p}\}_{p}$ and $\{\sigma^{x}_{j}\}_{j\in{\rm snake}}$. Here |$\ket{m}=\sigma^{x}_{j_{1}}\cdots\sigma^{x}_{j_{k}}\ket{+1^{L^{2}}}$ is defined as the bond configuration for $\{\textbf{Z}_{p}\}_{p}$ associated with the spin configuration $\sigma^{x}_{j_{1}}\cdots\sigma^{x}_{j_{k}}\ket{\psi_{o}}$, where |$\ket{\psi_{o}}$ is a ground state (4.5). This gives the alternative representation for |$\ket{m^{\prime}}\bra{m}$ |:
@@
Here, $\frac{I-m_{j_{1}}\textbf{Z}_{j_{1}}}{2}\ket{m^{\prime}}=\ket{m^{\prime}}$ and $\frac{I-m_{j_{1}}\textbf{Z}_{j_{1}}}{2}\ket{m}=0$. In the case when $j_{1}=1$, we ignore the condition $j\neq j_{1}-1$. The case of |$\ket{m}\bra{m}$ | can be similarly done. ∎